OpenStudy (anonymous):

Hello. Trying to exercise "simple" absolute value problems.. Here I got stuck again!

OpenStudy (anonymous):

\[\LARGE \begin{array}{l}6.\,\,\,\,\, - 3\left| { - 2x + 4} \right| \le 4\\\\\frac{{ - 3\left| { - 2x + 4} \right|}}{{ - 3}} \ge \frac{4}{{ - 3}}\\\left| { - 2x + 4} \right| \ge - \frac{4}{3}\\ - \left( { - \frac{4}{3}} \right) \ge - 2x + 4 \ge - \frac{4}{3}\\\frac{4}{3} - 4 \ge - 2x \ge - \frac{4}{3} - 4\\ - \frac{8}{3} \ge - 2x \ge - \frac{{16}}{3}\\\frac{8}{3} \le 2x \le \frac{{16}}{3}\\\frac{8}{{3 \cdot 2}} \le x \le \frac{{16}}{{3 \cdot 2}}\\\frac{8}{6} \le x \le \frac{{16}}{6}\\\frac{4}{3} \le x \le \frac{8}{3}\\x \in \left[ {\frac{4}{3},\frac{8}{3}} \right]\end{array}\] is this wrong or not... is it is, where... In the web it says that answer is (- infinity , +infinity) but I get a different answer :(

OpenStudy (anonymous):

I forgot to mentioned that this is not solved in web.. (steps) I did them :( ... Thanks

OpenStudy (anonymous):

yes... although I don't have |A| > |B| I have |A|>B but that it doesn't matter... from |A|>B -B>A>B is the same as A>B or A<-B and I think that's what I did. :(

OpenStudy (anonymous):

http://www.analyzemath.com/AbsEqIneq/answers_AbsEqIneq.html problem nr. 7

OpenStudy (anonymous):

Your inequality is true for every x. −3∣−2x+4∣<= 0 for all x so it is less than 4.

OpenStudy (anonymous):

Here's the intuitive explanation. −3*∣−2x+4∣ ≤4 That absolute value portion of the equation is always non-negative. Multiplying by a negative number will yield a negative, therefore the result will be less than 4. However, I'm not sure which part of your workout is wrong yet.

OpenStudy (anonymous):

but how.. why? ... Honestly I'm confused now :( −3∣−2x+4∣<= 0 ?? it's −3∣−2x+4∣<= 4 :) what's wrong with my answer.. :( I need some help here.

OpenStudy (anonymous):

The solution is I said before is \[ -\infty < x <\infty \]

OpenStudy (anonymous):

according to your solution, x=0 is not in there but it does work...

OpenStudy (anonymous):

@eliassaab is correct.

OpenStudy (anonymous):

Nevermind. I found the mistake you made. It was at this step. |-2x+4| >= -4/3 When it's a greater than sign, I can rewrite it as -2x+4 >= -4/3 OR -2x+4 <= -4/3 You used the rule for when it is a less than sign.

OpenStudy (anonymous):

@dpaInc I have no doubts that eliassaab might be wrong... but I want to know how this should be solved thenn... only by observation? :(

OpenStudy (anonymous):

\[-\frac{-4}{3} \ge -2x+4 \ge -\frac{4}{3}\] This step of your work was incorrect.

OpenStudy (anonymous):

Instead you should have \[-2x+4 \ge \frac{-4}{3} or -2x+4 \le \frac{-4}{3}\]

OpenStudy (anonymous):

@SmoothMath what you're saying it doesn't make sense now !! ... \[\LARGE -2x+4\geq -\frac43\] or \[\LARGE -2x+4\leq -\frac43\] is the same as: \[\LARGE -\frac43\geq -2x+4\geq -\frac43 \] which is exactly what I did...

OpenStudy (anonymous):